OpenStudy (anonymous):

the sum of three consecutive intergers is 93 find the produc of the intergers 29,760 32,736 35,904 24,360

OpenStudy (anonymous):

x+(x+1)+(x+2)=93 find x and then x+1 and x+2 then find their product.

OpenStudy (anonymous):

i don't understand???

OpenStudy (anonymous):

3x+3=93 x=? x+1=? x+2=?

OpenStudy (anonymous):

thank you mangorox

OpenStudy (anonymous):

but it doesn't show how to do a three consecutive integers

OpenStudy (anonymous):

3x=93-3=90 x=90/3=30 x+1=30+1=31 x+2=30+2=32 30*31*32=?

OpenStudy (anonymous):

but it the answers are different look above and i know its that but the answer are different ^^^